# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ Abstract
Motivated by the recent observations of $T^{*}_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ charmed tetraquark states by LHCb, we study the decays of heavy mesons to these charmed tetraquark states ($T$) using a topological amplitude approach. We first obtain the $T\to DP$ and $DS$ strong decay amplitudes by decomposing them into several topological amplitudes, where $P$ is a light psedo-scalar particle and $S$ is a low-lying scalar particle. Later on, weak decay amplitudes of $\overline{B}\to D\overline{T}$, $\overline{D}T$ and $\overline{B}\to TP$, $TS$ decays are also decomposed topologically. In addition, $B^{-}_{c}\to T\overline{T}$ decays are also discussed. Using these results, modes with an unambiguous exotic interpretation in flavor are highlighted.
@@ I Introduction
In 2020, LHCb reported the observation of a charmed tetraquark $T^{*}_{cs0}(2870)^{0}$ in $B^{-}\to D^{-}D^{+}K^{-}$ decay with $B^{-}\to D^{-}T^{*}_{cs0}(2870)^{0},T^{*}_{cs0}(2870)^{0}\to D^{+}K^{-}$ decay LHCb:2020pxc . From $T^{*}_{cs0}(2870)^{0}\to D^{+}K^{-}$ decay, the minimal quark content of $cs\bar{u}\bar{d}$ for $T^{*}_{cs0}(2870)^{0}$ is infered. Several years later, the observations of two other charmed tetraquarks, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$, are reported in $\overline{B}{}^{0}\to D^{0}D_{s}^{-}\pi^{+}$ and $B^{-}\to D^{+}D_{s}^{-}\pi^{-}$ decays with $\overline{B}{}^{0}\to D^{0}\overline{T^{*}_{c\bar{s}0}}(2900)^{0},\overline{T^{*}_{c\bar{s}0}}(2900)^{0}\to D_{s}^{-}\pi^{+}$ and $B^{-}\to D^{+}\overline{T^{*}_{c\bar{s}0}}(2900)^{\,--},\overline{T^{*}_{c\bar{s}0}}(2900)^{\,--}\to D_{s}^{-}\pi^{-}$ decays LHCb:2022sfr ; LHCb:2022lzp . From ${T^{*}_{c\bar{s}0}}(2900)^{0}\to D_{s}^{+}\pi^{-}$ and ${T^{*}_{c\bar{s}0}}(2900)^{\,++}\to D_{s}^{+}\pi^{+}$ decays the minimal quark contents of these states are $c\bar{s}\bar{u}d$ and $c\bar{s}u\bar{d}$, respectively. The masses and widths of these states are ParticleDataGroup:2024cfk
@@
There are many theoretical interpretations of these states; see, for example, refs. Karliner:2020vsi ; He:2020jna ; Wang:2020xyc ; Molina:2020hde ; Huang:2020ptc ; Liu:2020nil ; Wang:2021lwy ; Yue:2022mnf ; Burns:2020epm ; Liu:2020orv see also ref. PDGreview . In this work, we will study the decays of heavy mesons to these charmed tetraquark states ($T$). As a direct calculation of the decay amplitudes is too complicated, we will make use of a topological amplitude approach, which have been applied to many heavy meson decays, see, for example, refs. Zeppenfeld:1980ex ; Chau:tk ; Chau:1990ay ; Gronau:1994rj ; Gronau:1995hn ; Chiang:2004nm ; Cheng:2014rfa ; Savage:ub ; Li:2012cfa ; Qin:2013tje ; Chua:2003it ; Chua:2013zga ; Cheng:2015cca ; Chua:2016aqy ; Chua:2022wmr ; Chua:2023qyj . In particular, in ref. Cheng:2015cca , we applied this method to heavy baryons and pentaquark decays, obtaining results later confirmed experimentally LHCb:2016lve . Note that a study on tetraquark production in $B$ decays using diagrammatic analysis was given in ref. Qin:2022nof . We will compare our results with those obtained in that work later.
In this work, we will consider two scenarios where the light antiquarks in the tetraquarks are either antisymmetric or symmetric. It will be interesting to see which scenario the above tetraquarks, namely $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$, fit. We will consider the strong decays of tetraquarks and weak decays involving tetraquarks. Attentions are paid to predictions where all $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved. We will try to identify other flavor exotic states. It will be useful to identify modes involving flavor exotic states. We will be in a better position to understand these tetraquarks with this information.
The layout of this paper is as follows. We introduce two different scenarios of the light quark pair $\bar{q}\bar{q}$ in the charmed tetraquarks $T_{cq\bar{q}\bar{q}}$ in section II. All flavor exotic states are identified. We consider $T\to DP$ and $DS$ decays in Sec. III, $\overline{B}\to D\overline{T}$ and $\overline{D}T$ decays in Sec. IV, $\overline{B}\to TP$ and $TS$ decays in Sec. V and $B_{c}\to T\overline{T}$ decays in Sec. VI. Decay amplitudes of these decays are given in terms of topological amplitudes. These results can be easily generalized to modes with $D^{*}$ and $V$, instead of $D$ and $P$, respectively. Modes with unambiguous exotic interpretations in flavor are highlighted. Sec. VII comes to the conclusions. We end this paper with two appendices.
@@ II.1 Scenarios I and II
The tetraquark $T_{cq\bar{q}^{\prime}\bar{q}^{\prime\prime}}$ has two antiquarks. We can decompose the anti-quark pair $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ into anti-symmetric $[\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ and symmetric $\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}$ at the outset, since SU(3) transformation does not mix up these two sets. Hence, in this work, we consider two different configurations for the light anti-quarks in a $T_{cq\bar{q}^{\prime}\bar{q}^{\prime\prime}}$ state, namely an antisymmetric $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ configuration and a symmetric $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ configuration. They can be further separated into traceless and traceful parts, which do not mix under SU(3) transformation, see Appendix A.
@@
where the $[\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ notation indicates the antisymmetric configuration of these light anti-quarks. These nine ${\bf 3}\times{\bf 3}$ states can be grouped into a ${\bf 6}$ and a $\bar{\bf 3}$, see Eq. (133), where the ${\bf 6}$ is the traceless part, and it consists of an $S=+1$ isotriplet, $(T^{++}_{c\bar{s}},T^{+}_{c\bar{s}},T^{0}_{c\bar{s}})$, an $S=-1$ isosinglet, $T^{0}_{cs}$, and an $S=0$ isodoublet, $(T_{c}^{+},T_{c}^{0})$, while the $\bar{\bf 3}$ is the traceful part, and it has an $S=0$ isodoublet, $(T^{\prime\prime+}_{c},T^{\prime\prime 0}_{c})$, and an $S=+1$ isosinglet, $T^{\prime\prime+}_{c\bar{s}}$, see Fig. 1 (a). The quark contents and quantum numbers of these states are given in Table 1. Consequently, the above $3\times 3$ matrix ${\bf T}$ can be expressed as
@@
In this scenario, namely scenario I, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the ${\bf 6}$. In particular, $T^{*}_{c\bar{s}0}(2900)^{++}$ and $T_{c\bar{s}0}(2900)^{0}$ are $T^{++}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$, respectively, belonging to the $S=1$ isotriplet, while $T^{*}_{cs0}(2870)^{0}$ is $T^{0}_{cs}$ in the $S=-1$ isosinglet as shown in Table 1. The three flavor exotic states are labeled by asterisks in the table.
@@
where only the $\bar{\bf 3}$ states remain after taking the trace $T^{i}_{[ik]}$, as the ${\bf 6}$ is traceless, where one can also infer from Table 1.
Table 1: Quark contents and quantum numbers of charmed tetraquark in scenario I with antisymmetric $\bar{q}\bar{q}^{\prime}$. We have a ${\bf 6}$ and a $\bar{\bf 3}$. We numbered these states and states with flavor exotic are labeled by asterisks. In this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the ${\bf 6}$.
@@
$5$ $T^{+}_{c}$ $\frac{1}{2}(cu\bar{u}\bar{d}-cu\bar{d}\bar{u}+cs\bar{d}\bar{s}-cs\bar{s}\bar{d})$ $({\bf 6},\frac{1}{2},+\frac{1}{2},0)$
$6$ $T^{0}_{c}$ $\frac{1}{2}(cd\bar{u}\bar{d}-cd\bar{d}\bar{u}+cs\bar{s}\bar{u}-cs\bar{u}\bar{s})$ $({\bf 6},\frac{1}{2},-\frac{1}{2},0)$
$7$ $T^{\prime\prime+}_{c}$ $\frac{1}{2}(cu\bar{u}\bar{d}-cu\bar{d}\bar{u}-cs\bar{d}\bar{s}+cs\bar{s}\bar{d})$ $(\bar{\bf 3},\frac{1}{2},+\frac{1}{2},0)$
$8$ $T^{\prime\prime 0}_{c}$ $\frac{1}{2}(-cd\bar{u}\bar{d}+cd\bar{d}\bar{u}+cs\bar{s}\bar{u}-cs\bar{u}\bar{s})$ $(\bar{\bf 3},\frac{1}{2},-\frac{1}{2},0)$
$9$ $T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(-cu\bar{s}\bar{u}+cu\bar{u}\bar{s}+cd\bar{d}\bar{s}-cd\bar{s}\bar{d})$ $(\bar{\bf 3},0,0,+1)$
Table 2: Quark contents and quantum numbers of charmed tetraquark in scenario II with symmetric $\bar{q}\bar{q}^{\prime}$. We have a $\overline{\bf 15}$ and a $\bar{\bf 3}$. We numbered these states, and states with flavor exotic are labeled by asterisks. In this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\overline{\bf 15}$.
[…]
$1^{\prime*}$ $T^{++}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(cu\bar{d}\bar{s}+cu\bar{s}\bar{d})$ $(\overline{\bf 15},1,+1,+1)$ $T^{*}_{c\bar{s}0}(2900)^{++}$
@@
$3^{\prime*}$ $T^{0}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(cd\bar{s}\bar{u}+cd\bar{u}\bar{s})$ $(\overline{\bf 15},1,-1,+1)$ $T^{*}_{c\bar{s}0}(2900)^{0}$
$4^{\prime*}$ $T^{+}_{cs}$ $cs\bar{d}\bar{d}$ $(\overline{\bf 15},1,+1,-1)$
$5^{\prime*}$ $T^{0}_{cs}$ $\frac{1}{\sqrt{2}}(cs\bar{u}\bar{d}+cs\bar{d}\bar{u})$ $(\overline{\bf 15},1,0,-1)$ $T^{*}_{cs0}(2870)^{0}$
$6^{\prime*}$ $T^{-}_{cs}$ $cs\bar{u}\bar{u}$ $(\overline{\bf 15},1,-1,-1)$
$7^{\prime*}$ $T^{++}_{c\bar{s}\bar{s}}$ $cu\bar{s}\bar{s}$ $(\overline{\bf 15},\frac{1}{2},+\frac{1}{2},+2)$
$8^{\prime*}$ $T^{+}_{c\bar{s}\bar{s}}$ $cd\bar{s}\bar{s}$ $(\overline{\bf 15},\frac{1}{2},-\frac{1}{2},+2)$
$9^{\prime*}$ $T^{++}_{c}$ $cu\bar{d}\bar{d}$ $(\overline{\bf 15},\frac{3}{2},+\frac{3}{2},0)$
@@
$12^{\prime*}$ $T^{-}_{c}$ $cd\bar{u}\bar{u}$ $(\overline{\bf 15},\frac{3}{2},-\frac{3}{2},0)$
$13^{\prime}$ $T^{\prime+}_{c}$ $\frac{1}{2\sqrt{6}}(cu\bar{u}\bar{d}+cu\bar{d}\bar{d}+2cd\bar{d}\bar{d}-3cs\bar{d}\bar{s}-3cs\bar{s}\bar{d})$ $(\overline{\bf 15},\frac{1}{2},+\frac{1}{2},0)$
$14^{\prime}$ $T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{6}}(2cu\bar{u}\bar{u}+cd\bar{u}\bar{d}+cd\bar{d}\bar{d}-3cs\bar{s}\bar{u}-3cs\bar{u}\bar{s})$ $(\overline{\bf 15},\frac{1}{2},-\frac{1}{2},0)$
@@
$16^{\prime}$ $T^{\prime\prime+}_{c}$ $\frac{1}{2\sqrt{2}}(cu\bar{u}\bar{d}+cu\bar{d}\bar{d}+2cd\bar{d}\bar{d}+cs\bar{d}\bar{s}+cs\bar{s}\bar{d})$ $(\bar{\bf 3}^{\prime},\frac{1}{2},+\frac{1}{2},0)$
$17^{\prime}$ $T^{\prime\prime 0}_{c}$ $\frac{1}{2\sqrt{2}}(2cu\bar{u}\bar{u}+cd\bar{u}\bar{d}+cd\bar{d}\bar{d}+cs\bar{s}\bar{u}+cs\bar{u}\bar{s})$ $(\bar{\bf 3}^{\prime},\frac{1}{2},-\frac{1}{2},0)$

### v3

# Heavy Mesons to Open-Charmed Tetraquark Decays

@@ Abstract
Motivated by the recent observations of $T^{*}_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ open-charmed tetraquark states by LHCb, we study the decays of heavy mesons to these open-charmed tetraquark states ($T$) using a topological amplitude approach. We first obtain the $T\to DP$ and $DS$ strong decay amplitudes by decomposing them into several topological amplitudes, where $P$ is a light pseudo-scalar particle and $S$ is a low-lying scalar particle. Subsequently, weak decay amplitudes of $\overline{B}\to D\overline{T}$, $\overline{D}T$ and $\overline{B}\to TP$, $TS$ decays are decomposed topologically. In addition, $B^{-}_{c}\to T\overline{T}$ decays are also discussed. Using these results, modes with an unambiguous exotic interpretation in flavor are highlighted.
@@ I Introduction
In 2020, LHCb reported the observation of an open-charmed tetraquark $T^{*}_{cs0}(2870)^{0}$ in $B^{-}\to D^{-}D^{+}K^{-}$ decay with $B^{-}\to D^{-}T^{*}_{cs0}(2870)^{0},T^{*}_{cs0}(2870)^{0}\to D^{+}K^{-}$ decay LHCb:2020pxc . From $T^{*}_{cs0}(2870)^{0}\to D^{+}K^{-}$ decay, the minimal quark content of $cs\bar{u}\bar{d}$ for $T^{*}_{cs0}(2870)^{0}$ is infered. Several years later, the observations of two other open-charmed tetraquarks, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$, were reported in $\overline{B}{}^{0}\to D^{0}D_{s}^{-}\pi^{+}$ and $B^{-}\to D^{+}D_{s}^{-}\pi^{-}$ decays with $\overline{B}{}^{0}\to D^{0}\overline{T^{*}_{c\bar{s}0}}(2900)^{0},\overline{T^{*}_{c\bar{s}0}}(2900)^{0}\to D_{s}^{-}\pi^{+}$ and $B^{-}\to D^{+}\overline{T^{*}_{c\bar{s}0}}(2900)^{\,--},\overline{T^{*}_{c\bar{s}0}}(2900)^{\,--}\to D_{s}^{-}\pi^{-}$ decays LHCb:2022sfr ; LHCb:2022lzp . From ${T^{*}_{c\bar{s}0}}(2900)^{0}\to D_{s}^{+}\pi^{-}$ and ${T^{*}_{c\bar{s}0}}(2900)^{\,++}\to D_{s}^{+}\pi^{+}$ decays the minimal quark contents of these states are $c\bar{s}\bar{u}d$ and $c\bar{s}u\bar{d}$, respectively. The masses and widths of these states are ParticleDataGroup:2024cfk
@@
There are many theoretical interpretations of these states; see, for example, refs. Karliner:2020vsi ; He:2020jna ; Wang:2020xyc ; Molina:2020hde ; Huang:2020ptc ; Liu:2020nil ; Wang:2021lwy ; Yue:2022mnf ; Burns:2020epm ; Liu:2020orv , see also ref. PDGreview . In this work, we will study the decays of heavy mesons to these open-charmed tetraquark states ($T$). As a direct calculation of the decay amplitudes is too complicated, we will make use of a topological amplitude approach, which have been applied to many heavy meson decays, see, for example, refs. Zeppenfeld:1980ex ; Chau:tk ; Chau:1990ay ; Gronau:1994rj ; Gronau:1995hn ; Rosner:1999xd ; Cheng:2002ai ; Chiang:2004nm ; Cheng:2014rfa ; Savage:ub ; Li:2012cfa ; Qin:2013tje ; Chua:2003it ; Chua:2013zga ; Cheng:2015cca ; Chua:2016aqy ; Chua:2022wmr ; Chua:2023qyj . In particular, this approach can also be applied to decays involving exotic states. For example, this method was employed to study heavy baryons to pentaquark decays in ref. Cheng:2015cca , and the results obtained were confirmed by LHCb subsequently LHCb:2016lve . Note that a study on charmed tetraquark productions in $B$ decays using diagrammatic analysis was given in ref. Qin:2022nof . We will compare our results with those obtained in that work in later sections.
In this work, we will consider two scenarios where the light antiquarks in the open-charmed tetraquarks are either antisymmetric or symmetric. It will be interesting to see which scenario the above tetraquarks, namely $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$, fit. We will consider the strong decays of open-charmed tetraquarks and weak decays involving these tetraquarks. Attentions are paid to predictions where all $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved. We will try to identify other flavor exotic states as well. It will be useful to identify modes involving flavor exotic states as much as possible. We will be in a better position to understand these tetraquarks with this information.
The layout of this paper is as follows. We introduce two different scenarios of the light quark pair $\bar{q}\bar{q}$ in the open-charmed tetraquarks $T_{cq\bar{q}\bar{q}}$ in section II. All flavor exotic states are identified. We consider $T\to DP$ and $DS$ decays in Sec. III, $\overline{B}\to D\overline{T}$ and $\overline{D}T$ decays in Sec. IV, $\overline{B}\to TP$ and $TS$ decays in Sec. V and, finally, $B_{c}\to T\overline{T}$ decays in Sec. VI. Decay amplitudes of these decays are given in terms of topological amplitudes. These results can be easily generalized to modes with $D^{*}$ and $V$, instead of $D$ and $P$, respectively. Modes with unambiguous exotic interpretations in flavor are highlighted. Sec. VII comes to the conclusions. We end this paper with three appendices.
@@ II.1 Scenarios I and II
The open-charmed tetraquark $T_{cq\bar{q}^{\prime}\bar{q}^{\prime\prime}}$ has two light antiquarks, $\bar{q}^{\prime}\bar{q}^{\prime\prime}$. We can decompose the anti-quark pair $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ into anti-symmetric $\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ and symmetric $\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}$ at the outset, since SU(3) transformation does not mix up these two sets. Hence, in this work, we consider two different configurations for the light anti-quarks in a $T_{cq\bar{q}^{\prime}\bar{q}^{\prime\prime}}$ state, namely an antisymmetric $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ configuration and a symmetric $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ configuration. They can be further separated into traceless and traceful parts, which do not mix under SU(3) transformation, see Appendix A.
@@
where the $\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ notation indicates the antisymmetric configuration of these light anti-quarks. These nine ${\bf 3}\times{\bf 3}$ states can be grouped into a $\bf 6$} and a $\bar{\bf 3}$, see Eq. (175), where the $\bf 6$} is the traceless part, and it consists of an $S=+1$ isotriplet, $(T^{++}_{c\bar{s}},T^{+}_{c\bar{s}},T^{0}_{c\bar{s}})$, an S=1S=-1 isosinglet, $T^{0}_{cs}$, and an $S=0$ isodoublet, $(T_{c}^{+},T_{c}^{0})$; while the $\bar{\bf 3}$ is the traceful part, and it has an $S=0$ isodoublet, $(T^{\prime\prime+}_{c},T^{\prime\prime 0}_{c})$, and an $S=+1$ isosinglet, $T^{\prime\prime+}_{c\bar{s}}$, see Fig. 1 (a). The quark contents and quantum numbers of these states are given in Table 1. Consequently, the above $3\times 3$ matrix $\bf T$} can be re-expressed as
@@
Table 1: Quark contents and quantum numbers of open-charmed tetraquark in scenario I with antisymmetric $\bar{q}\bar{q}^{\prime}$. We have a $\bf 6$} and a $\bar{\bf 3}$. We numbered these states and states with flavor exotic are labeled by asterisks. In this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\bf 6$}.
[…]
55 $T^{+}_{c}$ $\frac{1}{2}(cu\bar{u}\bar{d}-cu\bar{d}\bar{u}+cs\bar{d}\bar{s}-cs\bar{s}\bar{d})$ $({\bf 6},\frac{1}{2},+\frac{1}{2},0)$
66 $T^{0}_{c}$ $\frac{1}{2}(cd\bar{u}\bar{d}-cd\bar{d}\bar{u}+cs\bar{s}\bar{u}-cs\bar{u}\bar{s})$ $({\bf 6},\frac{1}{2},-\frac{1}{2},0)$
77 $T^{\prime\prime+}_{c}$ $\frac{1}{2}(cu\bar{u}\bar{d}-cu\bar{d}\bar{u}-cs\bar{d}\bar{s}+cs\bar{s}\bar{d})$ $(\bar{\bf 3},\frac{1}{2},+\frac{1}{2},0)$
88 $T^{\prime\prime 0}_{c}$ $\frac{1}{2}(-cd\bar{u}\bar{d}+cd\bar{d}\bar{u}+cs\bar{s}\bar{u}-cs\bar{u}\bar{s})$ $(\bar{\bf 3},\frac{1}{2},-\frac{1}{2},0)$
99 $T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(-cu\bar{s}\bar{u}+cu\bar{u}\bar{s}+cd\bar{d}\bar{s}-cd\bar{s}\bar{d})$ $(\bar{\bf 3},0,0,+1)$
In this scenario, namely scenario I, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\bf 6$}. In particular, $T^{*}_{c\bar{s}0}(2900)^{++}$ and $T_{c\bar{s}0}(2900)^{0}$ are $T^{++}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$, respectively, belonging to the $S=1$ isotriplet, while $T^{*}_{cs0}(2870)^{0}$ is $T^{0}_{cs}$ in the S=1S=-1 isosinglet as shown in Table 1. The three flavor exotic states are labeled by asterisks in the table.
@@
where only the $\bar{\bf 3}$ states remain after taking the trace $T^{i}_{[ik]}$, as the $\bf 6$} is traceless, which can also be inferred from Table 1.
Table 2: Quark contents and quantum numbers of open-charmed tetraquark in scenario II with symmetric $\bar{q}\bar{q}^{\prime}$. We have a $\overline{\bf 15}$ and a $\bar{\bf 3}$. We numbered these states, and states with flavor exotic are labeled by asterisks. In this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\overline{\bf 15}$.
@@
$1^{\prime*}$ $T^{++}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(cu\bar{d}\bar{s}+cu\bar{s}\bar{d})$ $(\overline{\bf 15},1,+1,+1)$ $T^{*}_{c\bar{s}0}(2900)^{++}$ $T_{cu\bar{d}\bar{s}}$
@@
$3^{\prime*}$ $T^{0}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(cd\bar{s}\bar{u}+cd\bar{u}\bar{s})$ $(\overline{\bf 15},1,-1,+1)$ $T^{*}_{c\bar{s}0}(2900)^{0}$ $T_{cd\bar{u}\bar{s}}$
$4^{\prime*}$ $T^{+}_{cs}$ $cs\bar{d}\bar{d}$ $(\overline{\bf 15},1,+1,-1)$ $T_{cs\bar{d}\bar{d}}$
$5^{\prime*}$ $T^{0}_{cs}$ $\frac{1}{\sqrt{2}}(cs\bar{u}\bar{d}+cs\bar{d}\bar{u})$ $(\overline{\bf 15},1,0,-1)$ $T^{*}_{cs0}(2870)^{0}$ $T_{cs\bar{u}\bar{d}}$
$6^{\prime*}$ $T^{-}_{cs}$ $cs\bar{u}\bar{u}$ $(\overline{\bf 15},1,-1,-1)$ $T_{cs\bar{u}\bar{u}}$
$7^{\prime*}$ $T^{++}_{c\bar{s}\bar{s}}$ $cu\bar{s}\bar{s}$ $(\overline{\bf 15},\frac{1}{2},+\frac{1}{2},+2)$ $T_{cu\bar{s}\bar{s}}$
$8^{\prime*}$ $T^{+}_{c\bar{s}\bar{s}}$ $cd\bar{s}\bar{s}$ $(\overline{\bf 15},\frac{1}{2},-\frac{1}{2},+2)$ $T_{cd\bar{s}\bar{s}}$
$9^{\prime*}$ $T^{++}_{c}$ $cu\bar{d}\bar{d}$ $(\overline{\bf 15},\frac{3}{2},+\frac{3}{2},0)$ $T_{cu\bar{d}\bar{d}}$
@@
$12^{\prime*}$ $T^{-}_{c}$ $cd\bar{u}\bar{u}$ $(\overline{\bf 15},\frac{3}{2},-\frac{3}{2},0)$ $T_{cd\bar{u}\bar{u}}$
$13^{\prime}$ $T^{\prime+}_{c}$ $\frac{1}{2\sqrt{6}}(cu\bar{u}\bar{d}+cu\bar{d}\bar{u}+2cd\bar{d}\bar{d}-3cs\bar{d}\bar{s}-3cs\bar{s}\bar{d})$ $(\overline{\bf 15},\frac{1}{2},+\frac{1}{2},0)$
$14^{\prime}$ $T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{6}}(2cu\bar{u}\bar{u}+cd\bar{u}\bar{d}+cd\bar{d}\bar{u}-3cs\bar{s}\bar{u}-3cs\bar{u}\bar{s})$ $(\overline{\bf 15},\frac{1}{2},-\frac{1}{2},0)$
@@
$16^{\prime}$ $T^{\prime\prime+}_{c}$ $\frac{1}{2\sqrt{2}}(cu\bar{u}\bar{d}+cu\bar{d}\bar{u}+2cd\bar{d}\bar{d}+cs\bar{d}\bar{s}+cs\bar{s}\bar{d})$ $(\bar{\bf 3}^{\prime},\frac{1}{2},+\frac{1}{2},0)$
$17^{\prime}$ $T^{\prime\prime 0}_{c}$ $\frac{1}{2\sqrt{2}}(2cu\bar{u}\bar{u}+cd\bar{u}\bar{d}+cd\bar{d}\bar{u}+cs\bar{s}\bar{u}+cs\bar{u}\bar{s})$ $(\bar{\bf 3}^{\prime},\frac{1}{2},-\frac{1}{2},0)$